OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the second and third equations, then the first and second equations. 5x + 2y + z = -23x + 4y + 3z = 2-4x - 3y - 3z = 1 Answer 5x + 2y = -23x + 4y = 2 -x + y = 3-12x - 2y = 8 5x + 2y = -2-12x - 2y = 8 x - y = -33x + 4y = 2

OpenStudy (anonymous):

The answer I have is B........

OpenStudy (anonymous):

those are the equations ? 5x + 2y + z = -2 3x + 4y + 3z = 2 -4x - 3y - 3z = 1

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

in order to eliminate z from the second we multiply the first by 3 and then subtract the first from the second : 3x + 4y + 3z = 2 minus 15x +6y+3z= -6 = -12x -2y = 8

OpenStudy (anonymous):

It's either B, or c

OpenStudy (anonymous):

:)))

OpenStudy (anonymous):

oh wait they want the first and second ? why i saw third :O

OpenStudy (anonymous):

:( I find no encouragements in any of this.

OpenStudy (anonymous):

ok so second from third we just add the equations 3x + 4y + 3z = 2 + -4x - 3y - 3z = 1 and get -x +y = 3

OpenStudy (anonymous):

so bravo ! it is B:)

OpenStudy (anonymous):

Oh my gosh ;')

OpenStudy (anonymous):

I won!!!!!!!!!!!!!!!!!!!!! I understand!!!!!!!!!!!!!!!!!! :))))))))))Thank you!!!!!!!!!!!

OpenStudy (anonymous):

yw:)